Find the ratio of the electric and gravitational forces between two protons.

Grade:11

1 Answers

Aditi Chauhan
9 years ago
Sol. Given: mass of proton = 1.67 × 10^–27 kg = M base p k = 9 × 10^9 Charge of proton = 1.6 × 10^–19 c = C base p G = 6.67 × 10^–11 Let the separation be ‘r’ Fe = k(C base p)^2/r^2 , fg= F(M base p)^2/r^2 Now, Fe : Fg = K(C base p)^2/r^2 * r^2/G(M base p)^2 = 9 * 10^9 * (1.6 * 1`0^-19)^2/6.67 * 10^-11 * (1.67 * 10^-27)^2 = 9 * 2.56 * 10^38 = 1,24 * 10^38